### v2

# Funnel Synthesis via LMI Copositivity Conditions for Nonlinear Systems

Abstract
Funnel synthesis refers to a procedure for synthesizing a time-varying controlled invariant set and an associated control law around a nominal trajectory. The computation of the funnel involves solving a continuous-time differential equation or inequality, ensuring the invariance of the funnel. Previous approaches often compromise the invariance property of the funnel; for example, they may enforce the equation or the inequality only at discrete temporal nodes and do not have a formal guarantee of invariance at all times. This paper proposes a computational funnel synthesis method that can satisfy the invariance of the funnel without such compromises. We derive a finite number of linear matrix inequalities (LMIs) that imply the satifaction of a continuous-time differential linear matrix inequality guaranteeing the invariance of the funnel at all times from the initial to the final time. To this end, we utilize LMI conditions ensuring matrix copositivity, which then imply continuous-time invariance. The primary contribution of the paper is to prove that the resulting funnel is indeed invariant over a finite time horizon. We validate the proposed method via a three-dimensional trajectory planning and control problem with obstacle avoidance constraints.
keywords:
Controlled invariant funnel , Robust control , Differential linear matrix inequality
[…]
[label1]organization=Department of Aeronautics and Astronautics, University of Washington,addressline=3940 Benton Ln NE, city=Seattle, postcode=98195, state=WA, country=United States
1 Introduction
Funnels, also referred to as tubes in model predictive control (MPC) literature [1], represent a time-varying controlled invariant set around a nominal trajectory over a finite-time horizon. Funnel synthesis refers to a procedure not only computing the invariant funnel but also synthesizing an associated feedback control. The computed funnel with the associated controller can serve various purposes including robust motion planning and control [2] and trajectory planning with guaranteed feasibility [3].
To derive the invariance conditions of the funnel, Lyapunov theory, dissipative theory, or Hamilton-Jacobi (HJ) reachability analysis have been commonly employed [4, 3]. These approaches result in differential equations or inequalities that ensure the invariance of the funnel with respect to variables such as ellipsoidal parameters and feedback gains. For instance, when employing a quadratic Lyapunov (storage) function, the invariance condition derived by Lyapunov theory has the form of a differential linear matrix inequality (DLMI) [5]. One challenge is solving the resulting differential equation or DLMI over a finite horizon especially when state and input constraints are involved.
Various methods have been employed to compute the finite-horizon controlled invariant funnel for nonlinear systems [6, 4]. One notable approach is the use of sum-of-squares (SOS) programming [2]. This method has been applied to polynomial Lyapunov functions based on Lyapunov theory. One issue in the work is that the resulting invariance condition derived by Lyapunov theory is only enforced at discrete node points. The work in [7] extends this by ensuring continuous-time invariance by employing the linear interpolation in time and allowing discontinuities at each node point. They compute the forward reachable set of the system for an admissible input set sequentially, within each time interval between discrete node points. However, their approach is not applicable to our problem, where both the feedback control and the controlled-invariant funnel need to be computed in the presence of constraints. In [2, 7], to apply SOS programming, the system dynamics must be approximated as polynomials, yet these works neglect the error arising from such polynomial approximations and do not provide frameworks to account for this error. More importantly, these studies overlook state and input constraints except for the input saturation, and hence the objective of funnel computation merely focuses on minimizing the size to avoid constraint violations.
In [3], the computation of a quadratic funnel using a quadratic Lyapunov function is studied, directly imposing the state and input constraints. This approach transforms the resulting DLMI into a finite number of linear matrix inequalities (LMIs), assuming Jacobian matrices to be within known polytopes. However, this assumption generally does not hold when the Jacobian matrices are evaluated around arbitrary dynamically-feasible nominal trajectories. Our previous work [5] takes a different approach, computing the quadratic funnel through numerical optimal control using a multiple shooting method, without making the assumption on the Jacobian matrices. The downside of the previous work is that the resulting DLMI is only ensured at each temporal node, thereby risking the invariance condition violations between these nodes.
The proposed work synthesizes funnels based on uncertain linear time-varying (LTV) systems obtained via linearization of the original nonlinear dynamics around given nominal trajectories. Hence, the following papers on the analysis and synthesis of uncertain LTV systems is relevant to this paper: [8, 9, 10, 11]. The results in [9] provide a computational approach to analyze robust performance of uncertain LTV systems whose uncertainties are characterized by integral quadratic constraints (IQCs). This work is later extended in [8] to the synthesis of robust output-feedback controllers. The algorithm proposed in [8] alternately computes the variables describing the Lyapunov function for the closed-loop LTV system with a fixed feedback controller. Then it computes the feedback control gains while fixing the variables defining the underlying Lyapunov function until convergence. That is, they fix a set of solution variables while solving the rest, and then they solve for the solution variables that they fixed while the rest are now kept constant, and the process repeats in an iterative fashion. On the other hand, our proposed work characterizes nonlinearities using incremental quadratic constraints based on Lipschitz nonlinearities and constructs a single convex optimization problem whose solution provides both the funnel and feedback controller parameters.
This paper provides a computational funnel synthesis method for nonlinear systems under bounded disturbances over a finite-time horizon. The main contribution is proving that the computed funnel is indeed invariant over the entire time interval. To achieve this, we first derive an incremental dynamical system representing the behavior relative to the given nominal trajectory, modeled as an uncertain LTV system. The key idea is to approximate this LTV system to an uncertain linear parameter-varying (LPV) system where the approximation error is treated as a state-, input-, and disturbance-dependent uncertainty. This enables the derivation of a finite number of LMIs whose satisfaction guarantees the required continuous-time DLMI for the invariance of the funnel. In this derivation, we employ LMI conditions that ensure matrix copositivity [12, 13]. Including linear state and input constraints together, the funnel synthesis problem can be finally formulated as a finite-dimensional semidefinite program (SDP). Since both the system’s nonlinearity and LPV approximation errors are explicitly accounted for, the resulting funnel is guaranteed to be invariant for the original nonlinear system.
1.1 Notation
The set notations $\mathbb{R}$, $\mathbb{R}_{+}$, $\mathbb{Z}$ and $\mathbb{R}^{n}$ are the sets of real, nonnegative real, integer, and the $n$-dimensional Euclidean spaces, respectively. Intervals are written by $\mathbb{Z}_{[a,b)}=\{z\in\mathbb{Z}:a\leq z<b\}$. The symmetric matrix $Q=Q^{\top}(\succeq)\succ 0$ implies $Q$ is a positive (semi) definite matrix. The set of positive (semi) definite matrices whose size is $n\times n$ are denoted by $\mathbb{S}_{++}^{n}(\mathbb{S}_{+}^{n})$. The identity matrix having $n\times n$ size is denoted by $I_{n}$. The subscript and the time argument will be omitted when it is clear from the context. The notation $\mathcal{L}_{2}[a,b]$ is a set of Lebesgue measurable functions $x(t)$ defined on an interval $[a,b]\subset\mathbb{R}$ such that $\left(\int_{a}^{b}x(t)^{\top}x(t)\text{d}t\right)^{1/2}<\infty.$ A property is said to hold almost everywhere, or for almost every $x\in X$ with some set $X$ if the set on which it fails is a Lebesgue measure zero set. We abbreviate the notations $A^{\top}PA$ and $\left[\begin{array}[]{cc}a&b^{\top}\\ b&c\end{array}\right]$ as $(\star)^{\top}PA$ and $\left[\begin{array}[]{cc}a&\star\\ b&c\end{array}\right]$, respectively. The Minkowski sum is denoted by $\oplus$.
2 Invariance of funnel
2.1 Nonlinear systems
Consider the finite-horizon continuous-time nonlinear systems in the form of
$\dot{x}(t)=f(t,x(t),u(t),w(t)),\quad\forall t\in[t_{0},t_{f}],$ (1)
where $x(t)\in\mathbb{R}^{n_{x}}$ is the state, $u(t)\in\mathbb{R}^{n_{u}}$ is the input, $w(t)\in\mathbb{R}^{n_{w}}$ is the (exogenous) disturbance, and $t_{0},t_{f}$ are the initial and final time. The system dynamics $f:\mathbb{R}_{+}\times\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\times\mathbb{R}^{n_{w}}\rightarrow\mathbb{R}^{n_{x}}$ are assumed to be continuous in $t$ and continuously differentiable in $x,\,u$ and $w$. We assume that the input signal $u(\cdot)\in\mathcal{L}_{2}^{n_{u}}[t_{0},t_{f}]$ and the disturbance signal $w(\cdot)\in\mathcal{L}_{2}^{n_{w}}[t_{0},t_{f}]$ are piecewise continuous. We further assume that the disturbance signal $w(\cdot)$ is essentially bounded from above one, that is,
$\lVert w(\cdot)\rVert_{\infty}\leq 1,$ (2)
where $\lVert w(\cdot)\rVert_{\infty}\coloneqq\text{esssup}_{t\in[t_{0},t_{f}]}\lVert w(t)\rVert_{2}$. We refer to a trajectory as a collection of the state, the input, and the disturbance signals, denoted together by $(x(\cdot),u(\cdot),w(\cdot))$. The nominal trajectory having the zero disturbance $(\bar{x}(\cdot),\bar{u}(\cdot),0)$ is assumed to be given and dynamically feasible, that is, $\dot{\bar{x}}=f(t,\bar{x},\bar{u},0)$ for all $t\in[t_{0},t_{f}]$. With the given nominal trajectory, we can rewrite the system (1) in the linear fractional form
$f(t,x,u,w)$ $=A(t)x+B(t)u+F(t)w+E_{o}\phi(t,q)$ (3a)
$q$ $=C_{o}x+D_{o}u+G_{o}w,$ (3b)
where the function $\phi:\mathbb{R}_{+}\times\mathbb{R}^{n_{q}}\rightarrow\mathbb{R}^{n_{\phi}}$ represents the nonlinearity of the system and $q(t)\in\mathbb{R}^{n_{q}}$ is its argument. The matrices $A(t)$, $B(t)$, and $F(t)$ are chosen as first-order approximations of (1) around the nominal trajectory. We assume that $A(t)$, $B(t)$, and $F(t)$ are bounded for all $t\in[t_{0},t_{f}]$. The constant matrices $E_{o}\in\mathbb{R}^{n_{x}\times n_{\phi}}$, $C_{o}\in\mathbb{R}^{n_{q}\times n_{x}}$, $D_{o}\in\mathbb{R}^{n_{q}\times n_{u}}$, and $G_{o}\in\mathbb{R}^{n_{q}\times n_{w}}$ are selector matrices, composed of 0s and 1s, chosen to structure the nonlinearity of the system.
Example 1.
The unicycle model can be written as
$\dot{x}=\left[\begin{array}[]{c}\dot{x}_{1}\\ \dot{x}_{2}\\ \dot{x}_{3}\end{array}\right]=\left[\begin{array}[]{c}u_{1}\cos(x_{3}+c_{1}w_{1})\\ u_{1}\sin(x_{3}+c_{1}w_{1})\\ u_{2}+c_{2}w_{2}\end{array}\right],$ (4)
where $x_{1}$ and $x_{2}$ are $x$- and $y$-positions, $x_{3}$ are the yaw angle, $u_{1}$ is the velocity control, and $u_{2}$ is the angular velocity control. The disturbances $w_{1}$ and $w_{2}$ affect to the yaw angle estimation and the angular velocity control, and $c_{1},c_{2}\in\mathbb{R}_{+}$ are system parameters. The argument $q$ of $\phi$ can be chosen as $[x_{3},u_{1},w_{1}]$ with
$C_{o}$ $=\left[\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 0&0&0\end{array}\right],D_{o}=\left[\begin{array}[]{cc}0&0\\ 1&0\\ 0&0\end{array}\right],G_{o}=\left[\begin{array}[]{cc}0&0\\ 0&0\\ 1&0\end{array}\right].$
Since the only first two components in $f$ involve with the nonlinearity, the matrix $E_{o}$ is $\left[\begin{array}[]{ccc}1&0&0\\ 0&1&0\end{array}\right]^{\top}$. The time-varying matrices $A(t),\,B(t)$, and $F(t)$ are given as follows:
$A(t)$ $=\left[\begin{array}[]{ccc}0&0&-\bar{u}_{1}(t)\sin\bar{x}_{3}(t)\\ 0&0&\bar{u}_{1}(t)\cos\bar{x}_{3}(t)\\ 0&0&0\end{array}\right],B(t)=\left[\begin{array}[]{cc}\cos\bar{x}_{3}(t)&0\\ \sin\bar{x}_{3}(t)&0\\ 0&1\end{array}\right],$ (5g)
$F(t)$ $=\left[\begin{array}[]{cc}-c_{1}\bar{u}_{1}(t)\sin\bar{x}_{3}(t)&0\\ c_{1}\bar{u}_{1}(t)\cos\bar{x}_{3}(t)&0\\ 0&c_{2}\end{array}\right]$ (5k)
where $\bar{x}=[\bar{x}_{1},\bar{x}_{2},\bar{x}_{3}]^{\top}$ adn $\bar{u}=[\bar{u}_{1},\bar{u}_{2}]^{\top}$ are the nominal state and input, respectively. The nonlinear function $\phi$ that is the remainder term in (3) is given by
$\phi(t,q)=\left[\begin{array}[]{c}c_{1}\bar{u}_{1}w_{1}\sin\bar{x}_{3}-u_{1}\cos\bar{x}_{3}+u_{1}\cos(c_{1}w_{1}+x_{3})+\bar{u}_{1}x_{3}\sin\bar{x}_{3},\\ -c_{1}\bar{u}_{1}w_{1}\cos\bar{x}_{3}-u_{1}\sin\bar{x}_{3}+u_{1}\sin(c_{1}w_{1}+x_{3})-\bar{u}_{1}x_{3}\cos\bar{x}_{3}\end{array}\right],$
where the time argument is omitted.
2.2 Incremental dynamical system and its LPV approximation
The incremental form of dynamics [14, 15, 16] illustrates the behavior of the system (1) relative to the nominal trajectory. To derive it, we first define difference variables as
$\eta(t)\coloneqq x(t)-\bar{x}(t),\quad\xi(t)=u(t)-\bar{u}(t),\quad\delta q(t)=q(t)-\bar{q}(t),$
where $\bar{q}(t)\coloneqq C_{o}\bar{x}(t)+D_{o}\bar{u}(t)$. Having $\eta$ as the state, the incremental dynamics can be derived as the following uncertain LTV system:
$\dot{\eta}$ $=A(t)\eta+B(t)\xi+F(t)w+E_{o}\left(\phi(t,q)-\phi(t,\bar{q})\right),$ (6a)
$\delta q$ $=C_{o}\eta+D_{o}\xi+G_{o}w.$ (6b)
Since the original system $f$ in (1) is continuously differentiable, $f$ is locally Lipschitz, thereby resulting in $\phi$ being locally Lipschitz with its second argument. Then, for each $t\in[t_{0},t_{f}]$, there exists a local Lipschitz constant $\gamma_{i}(t)\in\mathbb{R}_{+}$ such that
$\lVert\phi_{i}(t,q)-\phi_{i}(t,\bar{q})\rVert_{2}\leq\gamma_{i}(t)\lVert q(t)-\bar{q}(t)\rVert_{2},\quad\forall i\in\mathbb{Z}_{[1,n_{\phi}]},\quad\forall q,\bar{q}\in\mathcal{Q},$ (7)
for any compact set $\mathcal{Q}$. Here the term $\phi(t,q)-\phi(t,\bar{q})$, due to the nonlinearity of (3), is a state-, input-, and disturbance-dependent uncertainty characterized by the constraint (7).
Now we describe an LPV approximation of the LTV system (6). The underlying motivation of this approximation is that we can derive a finite number of LMIs that imply the invariance of the funnel for the LPV system, which is not tractable for the LTV system. We start by partitioning the time horizon $[t_{0},t_{f}]$ into $N\in\mathbb{R}_{+}$ uniform subintervals using time nodes defined as:
$t_{k}=t_{0}+\frac{k}{N}(t_{f}-t_{0}),\quad\forall k\in\mathbb{Z}_{[0,N]}.$
On each subinterval $[t_{k},t_{k+1}]$, we approximate the time-varying system matrices $A(t)$, $B(t)$, and $F(t)$ using a first-order hold (FOH) approach, which linearly interpolates between their values at the endpoints of the interval. We define:
~$\tilde{\square}(t)$ $\coloneqq\sigma_{1}^{k}(t)\square_{k}+\sigma_{2}^{k}(t)\square_{k+1},\quad t\in[t_{k},t_{k+1}]$ (8a)
$\sigma_{1}^{k}(t)$ $=\frac{t_{k+1}-t}{t_{k+1}-t_{k}},\quad\sigma_{2}^{k}(t)=\frac{t-t_{k}}{t_{k+1}-t_{k}},$ (8b)
where $\square_{k}=\square(t_{k})$ and the placeholder $\square$ corresponds to $A$, $B$, and $F$. By applying (8) across all $k$ in $\mathbb{Z}_{[0,N-1]}$, we obtain continuous piecewise linear approximations $(\tilde{A}(t),\tilde{B}(t),\tilde{F}(t))$ of the system matrices $(A(t),B(t),F(t))$ over the entire horizon $[t_{0},t_{f}]$.
Using (8), we can rewrite (6) equivalently as follows:
$\dot{\eta}$ $=\tilde{A}(t)\eta+\tilde{B}(t)\xi+\tilde{F}(t)w+\underbrace{\Delta(t)\left[\begin{array}[]{c}\eta\\ \xi\\ w\end{array}\right]}_{\coloneqq e(t)}+E_{o}\delta\phi,$ (9d)
$\Delta(t)$ $\coloneqq\left[\begin{array}[]{ccc}A(t)-\tilde{A}(t)&B(t)-\tilde{B}(t)&F(t)-\tilde{F}(t)\end{array}\right],$ (9f)
where $\delta\phi(t,\delta q)\coloneqq\phi(t,q)-\phi(t,\bar{q})$. Note that two representations (6) and (9) are equivalent by the definition of the error term $e(t)$ in (9d). Since $f$ is assumed to be continuously differentiable, the matrices $A(t)$, $B(t)$, and $F(t)$ are continuous, so bounded within the finite interval $[t_{0},t_{f}]$. Hence, there exists a positive constant $\beta(t)\in\mathbb{R}_{+}$ for each $t$ such that
$\lVert\Delta(t)\rVert_{2}\leq\beta(t),$ (10)
where $\lVert\cdot\rVert_{2}$ denotes the matrix 2-norm.
Alternatively, we can express the error term $e(t)$ more compactly as
$e(t)=E_{\Delta}\Delta(t)q_{\Delta}(t)$
where the matrix $E_{\Delta}\in\mathbb{R}^{n_{x}\times n_{\Delta}}$, the block-diagonal matrix $\Delta\in\mathbb{R}^{n_{\Delta}\times n_{\Delta}}$ of uncertainties, and the vector $q_{\Delta}\in\mathbb{R}^{n_{\Delta}}$ are given by
$E_{\Delta}$ $=\left[\begin{array}[]{ccc}E_{\eta}&E_{\xi}&E_{w}\end{array}\right],\quad\Delta=\left[\begin{array}[]{ccc}\Delta_{\eta}&0&0\\ 0&\Delta_{\xi}&0\\ 0&0&\Delta_{w}\end{array}\right],\quad q_{\Delta}=\left[\begin{array}[]{c}q_{\eta}\\ q_{\xi}\\ q_{w}\end{array}\right],$
with $n_{\Delta}=n_{q_{\eta}}+n_{q_{\xi}}+n_{q_{w}}$. These components are chosen such that
$(A(t)-\tilde{A}(t))\eta$ $=E_{\eta}\Delta_{\eta}q_{\eta},$
$(B(t)-\tilde{B}(t))\xi$ $=E_{\xi}\Delta_{\xi}q_{\xi},$
$(F(t)-\tilde{F}(t))w$ $=E_{w}\Delta_{w}q_{w},$
where the vectors $q_{\eta}$, $q_{\xi}$, and $q_{w}$ are linear functions of $\eta$, $\xi$, and $w$, respectively, as follows:
$q_{\eta}=C_{\eta}\eta,\quad q_{\xi}=D_{\xi}\xi,\quad q_{w}G_{w}w.$
In this formulation, the uncertainties are captured using component-wise inequalities:
$|\Delta_{i}(t)|\leq\beta_{i}(t),\quad i=\mathbb{Z}_{[1,n_{\Delta}]},$ (11)
where $\beta_{i}(t)>0$ bounds the corresponding uncertainty $\Delta_{i}(t)$. Compared to the single norm-bounded inequality in (10), these component-wise bounds in (11) can represent the uncertainly more compactly by exploiting its structure. More details on uncertainty representation can be found in [17].
Now we can summarize the incremental dynamics, written in the form of an uncertain LPV system, as follows:
$\dot{\eta}$ $=\tilde{A}(t)\eta+\tilde{B}(t)\xi+\tilde{F}(t)w+Ep(t,r),$ (12a)
$r$ $\coloneqq C\eta+D\xi+Gw=[q_{\Delta}^{\top},\delta q^{\top}]^{\top}$ (12b)
$p$ $\coloneqq\left[\begin{array}[]{c}\Delta(t)q_{\Delta}(t)\\ \delta\phi(t,\delta q)\end{array}\right],\quad E\coloneqq\left[\begin{array}[]{cc}E_{\Delta}&E_{o}\end{array}\right],$ (12f)
$C$ $\coloneqq\left[\begin{array}[]{c}C_{\eta}\\ 0\\ 0\\ C_{o}\end{array}\right],D\coloneqq\left[\begin{array}[]{c}0\\ D_{\xi}\\ 0\\ D_{o}\end{array}\right],G\coloneqq\left[\begin{array}[]{c}0\\ 0\\ G_{w}\\ G_{o}\end{array}\right].$ (12s)
All uncertainties in the system (12) are lumped into the vector $p(t,r)\in\mathbb{R}^{n_{p}}$ where $n_{p}=n_{\Delta}+n_{\phi}$ and $r(t)\in\mathbb{R}^{n_{r}}$. We characterize the uncertain term $p$ using the following quadratic inequality (QI) [18, 16] for each $t$:
$\left[\begin{array}[]{c}r(t)\\ p(t,r)\end{array}\right]^{\top}M(t)\left[\begin{array}[]{c}r(t)\\ p(t,r)\end{array}\right]\geq 0,$ (13)
where the matrix $M(t)\in\mathbb{R}^{(n_{r}+n_{p})\times(n_{r}+n_{p})}$ is called a multiplier matrix. It follows from (7) and (10) that the valid multiplier matrix has the form of
$M(t)$ $=\left[\begin{array}[]{cc}N_{1}(t)^{-1}&0\\ 0&-N_{2}(t)^{-1}\end{array}\right],N_{1}(t)\coloneqq\left[\begin{array}[]{cc}\lambda_{\beta}(t)I_{n_{r}-n_{q}}&0\\ 0&\lambda_{\gamma}(t)I_{n_{q}}\end{array}\right],$ (14e)
$N_{2}(t)$ $\coloneqq\left[\begin{array}[]{cc}\beta(t)^{2}\lambda_{\beta}(t)I_{n_{\Delta}}&0\\ 0&\lambda_{\gamma}(t)\Gamma(t)^{2}\end{array}\right],$ (14h)
where $\Gamma(t)=\text{diag}(\gamma_{1}(t),\ldots,\gamma_{n_{\phi}}(t))$ and any positive real-valued functions $\lambda_{\beta}(t)>0$ and $\lambda_{\gamma}(t)>0$. This can be verified by deriving the following:
$\left[\begin{array}[]{c}r\\ p\end{array}\right]^{\top}M\left[\begin{array}[]{c}r\\ p\end{array}\right]$ $=\frac{1}{\lambda^{\beta}}\left(\beta^{2}(\eta^{\top}\eta+\xi^{\top}\xi+w^{\top}w)-e^{\top}e\right)$
$+\frac{1}{\lambda^{\gamma}}\left(\delta q^{\top}\Gamma(t)^{2}\delta q-\delta\phi^{\top}\delta\phi\right)\geq 0,$
where the last inequality comes from (7) and (10). The resulting interconnection in the uncertain LPV system is illustrated in Fig. 1.
Uncertainties satisfying (13) LPV System (12a)Feedback Control (17)$p$$r$$w$$\eta$$\xi$$\eta$
Figure 1: Uncertain LPV system interconnection with feedback control.
2.3 Invariance condition
We consider a scalar-valued continuous Lyapunov function $V:\mathbb{R}\times\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$ defined by
$V(t,\eta(t))=\eta(t)^{\top}Q(t)^{-1}\eta(t),$ (15)
where $Q(t)\in\mathbb{S}_{++}^{n_{x}}$ is a continuous-time PD matrix-valued continuous function. A state funnel is defined as a 1-sublevel set of $V$, that is,
$\mathcal{E}_{Q}(t)\coloneqq\{\eta\mid\eta^{\top}Q(t)^{-1}\eta\leq 1\}.$ (16)
We employ the following linear time-varying feedback controller:
$\xi(t)=K(t)\eta(t),$ (17)
for the system (12) with the feedback gain $K(t)\in\mathbb{R}^{n_{u}\times n_{x}}$. With this linear feedback, $\xi$ becomes linear in $\eta$. It follows that $\eta\in\mathcal{E}_{Q}$ implies $\xi\in\mathcal{E}_{KQK^{\top}}$ [17] where
$\mathcal{E}_{KQK^{\top}}(t)\coloneqq\{(K(t)Q(t)K(t)^{\top})^{\frac{1}{2}}y\mid\lVert y\rVert_{2}\leq 1,y\in\mathbb{R}^{n_{u}}\}.$ (18)
We refer $\mathcal{E}_{KQK^{\top}}$ to as an input funnel. Notice that if $KQK^{\top}$ is PD, the set $\mathcal{E}_{KQK^{\top}}$ is equivalent to $\{\xi\mid\xi^{\top}(KQK^{\top})^{-1}\xi\leq 1\}$. Using the state and input funnels, a funnel $\mathcal{F}(t)$ is defined as an Cartesian product of state and input funnel centered around the nominal state $\bar{x}(t)$ and input $\bar{u}(t)$, respectively, that can be written as
$\mathcal{F}(t)\coloneqq\left(\{\bar{x}(t)\}\oplus\mathcal{E}_{Q}(t)\right)\times\left(\{\bar{u}(t)\}\oplus\mathcal{E}_{KQK^{\top}}(t)\right).$ (19)
Lemma 1.
Suppose there exists a PD matrix-valued continuous function $Q(t)\in\mathbb{S}_{++}^{n_{x}}$ and a continuous matrix-valued function $K(t)\in\mathbb{R}^{n_{u}\times n_{x}}$ such that all trajectories of the system (12) satisfy
$\text{If }V\geq w^{\top}w,\ \text{then }\dot{V}(t,\eta(t))\leq 0,$ (20)
for almost all $t\in[t_{0},t_{f}]$. Then, with every almost everywhere continuous signal $w(\cdot)$ such that (2), $\mathcal{E}_{Q}(t)$ defined in (16) is an invariant time-varying ellipsoid for the system (12) on $[t_{0},t_{f}]$, that is, if $\eta(\cdot)$ is a solution of (12) with $\eta(t_{0})\in\mathcal{E}_{Q}(t_{0})$, then $\eta(t)\in\mathcal{E}_{Q}(t)$ for all $t\in[t_{0},t_{f}]$.
Proof.
We prove by contradiction. Suppose that the invariance condition does not hold. Then, there exists a solution $\eta(\cdot)$ of (12) such that $\eta(t_{0})\in\mathcal{E}_{Q}(t_{0})$ and $\eta(t_{2})\notin\mathcal{E}_{Q}(t_{2})$ for some $t_{2}>t_{0}$. Since $V(t,\eta)$ is continuous in $t$ by continuity of $\eta$ and $Q$, there exists $t_{1}\in(t_{0},t_{2})$ such that $V(t_{1},\eta(t_{1}))=1$ and $V(t,\eta(t))>1$ for all $t\in(t_{1},t_{2}]$ by the intermediate value theorem. It follows from (20) that $\dot{V}(t,\eta(t))\leq 0$ for almost all $t\in[t_{1},t_{2}]$ as (2). Observe that
$V(t_{2},\eta(t_{2}))$ $=V(t_{1},\eta(t_{1}))+\int_{t_{1}}^{t_{2}}\dot{V}(t,\eta(t))\text{d}t,$
$\leq V(t_{1},\eta(t_{1}))=1.$
This implies $\eta(t_{2})\in\mathcal{E}_{Q}(t_{2})$ that contradicts our hypothesis that $\eta(t_{2})\notin\mathcal{E}_{Q}(t_{2})$. Thus, $\mathcal{E}_{Q}(t)$ is invariant. ∎
While Lemma 1 follows a similar proof as in [19, Lemma 1], it extends the result to consider more general types of disturbance signals $w(\cdot)$ that are continuous almost everywhere. In the next lemma, we define the invariance of the funnel $\mathcal{F}$ and its relevance to the invariance of $\mathcal{E}_{Q}$.
Lemma 2.
If $\mathcal{E}_{Q}(t)$ is invariant for the uncertain LPV system (12) with the linear feedback control $\xi=K\eta$, then the funnel $\mathcal{F}(t)$ defined in (19) is invariant for the original system (1) in a sense that if $x(\cdot)$ is a solution of (1) with $x(t_{0})\in\{\bar{x}(t)\}\oplus\mathcal{E}_{Q}(t)$ and the control law,
$u(t)=\bar{u}(t)+K(t)(x(t)-\bar{x}(t)),$ (21)
then $(x(t),u(t))\in\mathcal{F}(t)$ for all $t\in[t_{0},t_{f}]$.
Proof.
Consider a solution $x(\cdot)$ of (1) with $x(t_{0})\in\{\bar{x}(t_{0})\}\oplus\mathcal{E}_{Q}(t_{0})$. This implies that $\eta(t_{0})\in\mathcal{E}_{Q}(t_{0})$, and then by the invariance of $\mathcal{E}_{Q}$, we have $\eta(t)=x(t)-\bar{x}(t)\in\mathcal{E}_{Q}(t)$ with $\xi(t)=u(t)-\bar{u}(t)=K(t)\eta(t)$. Thus, $(x(t),u(t))\in\mathcal{F}(t)$ for all $t\in[t_{0},t_{f}]$ with (21). ∎
2.4 Differential linear matrix inequality
This subsection describes the derivation of the DLMI that implies the invariance condition (20).
Lemma 3.
Consider the uncertain LPV system in (12) with the multiplier matrix $M(t)$ satisfying (14). Suppose there exist $Q(t)\in\mathbb{S}_{++}^{n_{x}}$, $Y(t)\in\mathbb{R}^{n_{u}\times n_{x}}$, $\lambda_{\beta}(t)\in\mathbb{R}_{+}$, $\lambda_{\gamma}(t)\in\mathbb{R}_{+}$, and $\lambda_{w}\in\mathbb{R}_{+}$ such that the following differential matrix inequality holds for some $t\in[t_{0},t_{f}]$:
$D_{LMI}(t,\dot{Q},Q,Y,\lambda_{\beta},\lambda_{\gamma})$$\coloneqq$
$\left[\begin{array}[]{cccc}W-\dot{Q}&\star&\star&\star\\ N_{2}E^{\top}&-N_{2}&\star&\star\\ \tilde{F}^{\top}&0&-\lambda_{w}I&\star\\ CQ+DY&0&G&-N_{1}\end{array}\right]$ $\preceq 0,$ (26)
where $W=\tilde{A}Q+\tilde{B}Y+Q\tilde{A}^{\top}+Y^{\top}\tilde{B}^{\top}+\lambda_{w}Q$ and $N_{1},\,N_{2}$ are defined in (14). Let $K(t)=Y(t)Q(t)^{-1}$, and then the condition (20) holds for the given $t$.
Proof.
Pre- and post-multiplying (26) by $\text{diag}(Q^{-1},N_{2}^{-1},I,I)$ generates
$\left[\begin{array}[]{cccc}Q^{-1}(W-\dot{Q})Q^{-1}&\star&\star&\star\\ E^{\top}Q^{-1}&-N_{2}^{-1}&\star&\star\\ \tilde{F}^{\top}Q^{-1}&0&-\lambda_{w}I&\star\\ C_{cl}&0&G&-N_{1}\end{array}\right]\preceq 0,$
where $C_{cl}=C+DK$. By applying Schur complement, we can derive
$\left[\begin{array}[]{ccc}Q^{-1}(W-\dot{Q})Q^{-1}&\star&\star\\ E^{\top}Q^{-1}&0&\star\\ \tilde{F}^{\top}Q^{-1}&0&-\lambda_{w}I\end{array}\right]+(\star)^{\top}\left[\begin{array}[]{cc}N_{1}^{-1}&0\\ 0&N_{2}^{-1}\end{array}\right]\left[\begin{array}[]{ccc}C_{cl}&0&G\\ 0&I&0\end{array}\right]\preceq 0.$
It follows that
$\dot{V}+(\star)^{\top}\left[\begin{array}[]{cc}N_{1}^{-1}&0\\ 0&N_{2}^{-1}\end{array}\right]\left[\begin{array}[]{c}r\\ p\end{array}\right]+\lambda_{w}(V-w^{\top}w)\leq 0,$
for all $\eta\in\mathbb{R}^{n_{x}}$, $p\in\mathbb{R}^{n_{p}}$, and $w\in\mathbb{R}^{n_{w}}$. By using (13), S-procedure [17], and $\lambda_{w}>0$, we can conclude that the above inequality implies (20). ∎
We omit the time arguments in Lemma 3 and its proof for the notational brevity. The matrices $C,\,D,\,E,\,G$, and the constant $\lambda_{w}$ are time-invariant, while the terms $\tilde{A},\,\tilde{B},\,\tilde{F},\,Q,\,Y,\,K,\,W,\,N_{1}$, and $N_{2}$ are time-varying. Notice that with the fixed positive constant $\lambda_{w}$, $D_{LMI}(\dot{Q},Q,Y,\lambda_{\beta},\lambda_{\gamma})$ in (26) is linear in its arguments.
3 Funnel synthesis problem
3.1 Transformation of DLMI into a finite number of LMIs
To transform the DLMI (26) into a finite number of LMIs, we first parameterize $Q(t)$, $Y(t)$, $\lambda_{\beta}(t)$, and $\lambda_{\gamma}(t)$ by emplyoing the same FOH interpolation used in (8), that is,
~$\tilde{\square}(t)$ $\coloneqq\sigma_{1}^{k}(t)\square_{k}+\sigma_{2}^{k}(t)\square_{k+1},\quad t\in[t_{k},t_{k+1}],$
where a placeholder $\square$ corresponds to $Q,\,Y,\,\lambda_{\beta}$, and $\lambda_{\gamma}$, and the time-varying parameters $\sigma_{1}^{k}(t)$ and $\sigma_{2}^{k}(t)$ are given in (8b). Here $Q_{k}\in\mathbb{S}_{++}^{n_{x}},Y_{k}\in\mathbb{R}^{n_{u}\times n_{x}}$, $(\lambda_{\beta})_{k}\in\mathbb{R}_{+}$, and $(\lambda_{\gamma})_{k}\in\mathbb{R}_{+}$ for all $k\in\mathbb{Z}_{[0,N]}$ are our decision variables to be optimized. With this interpolation, the derivative of $\dot{Q}$ for each open subinterval $(t_{k},t_{k+1})$, denoted by $\dot{Q}^{k}$, is constant and given by
$\dot{Q}^{k}\coloneqq\dot{Q}=\frac{Q_{k+1}-Q_{k}}{t_{k+1}-t_{k}},$ (27)
where the superscript $k$ indicates that $\dot{Q}$ is specific to each subinterval.
To derive a finite number of LMIs that ensure the invariance, we also need to represent the time-varying constants $\gamma(t)$ and $\beta(t)$ with a finite set of values. To this end, we define
$(\gamma_{i})_{k}\coloneqq\sup_{t\in[t_{k},t_{k+1}]}\gamma_{i}(t),\quad\beta_{k}\coloneqq\sup_{t\in[t_{k},t_{k+1}]}\beta(t),$ (28)
for all $k$ in $\mathbb{Z}_{[0,N-1]}$. These constants $(\gamma_{i})_{k}$ and $\beta_{k}$ serve as upper bounds for $\gamma_{i}(t)$ and $\beta(t)$, respectively, on each subinterval $[t_{k},t_{k+1}]$. Hence, they become valid constants for (7) and (10), respectively on each subinterval $[t_{k},t_{k+1}]$. This allows us to incorporate $(\gamma_{i})_{k}$ and $\beta_{k}$ as constants in the LMIs associated with each subinterval, facilitating the derivation of a finite set of LMIs for the entire horizon $[t_{0},t_{f}]$
With the FOH interpolation (8) for the system matrices and our decision variables, the DLMI condition (26) for open subinterval $(t_{k},t_{k+1})$ can be equivalently rewritten as
$\sigma_{1}^{k}(t)\sigma_{1}^{k}(t)H_{k,k}^{k}+\sigma_{2}^{k}(t)\sigma_{2}^{k}(t)H_{k+1,k+1}^{k}+\sigma_{1}^{k}(t)\sigma_{2}^{k}(t)(H_{k,k+1}^{k}+H_{k+1,k}^{k})\succeq 0,$ (29)
where $H_{i,j}^{k}\in\mathbb{R}^{n_{H}\times n_{H}}$ with $n_{H}=n_{x}+n_{p}+n_{w}+n_{r}$ for $i,j\in\{k,k+1\}$ is given by
$H_{i,j}$ $\coloneqq-\left[\begin{array}[]{cccc}W_{i,j}-\dot{Q^{k}}&\star&\star&\star\\ N_{2,j}^{k}E^{\top}&-N_{2,j}^{k}&\star&\star\\ F_{j}^{\top}&0&-\lambda_{w}I&\star\\ L_{j}&0&G&-N_{1,j}\end{array}\right],$ (34)
$W_{i,j}$ $\coloneqq A_{i}Q_{j}+B_{i}Y_{j}+Q_{j}A_{i}^{\top}+Y_{j}^{\top}B_{i}^{\top}+\lambda_{w}Q_{j},$
$L_{j}$ $\coloneqq CQ_{j}+DY_{j},N_{1,j}\coloneqq\text{diag}((\lambda_{\beta})_{j}I,(\lambda_{\gamma})_{j}I),$
$N_{2,j}^{k}$ $\coloneqq\text{diag}(\beta_{j}^{2}(\lambda_{\beta})_{j}I,(\lambda_{\gamma})_{j}\Gamma_{k}^{2}),$
where $\Gamma_{k}=\text{diag}((\gamma_{1})_{k},\ldots,(\gamma_{n_{\phi}})_{k})$ and the superscript $k$ is introduced for the piecewise constant terms $\dot{Q}^{k}$, $\Gamma_{k}$, and $\beta_{k}$. Then, observe that (29) can be expressed as
$(\star)^{\top}\underbrace{\left[\begin{array}[]{cc}H_{k,k}^{k}&\star\\ \frac{1}{2}\left(H_{k,k+1}^{k}+H_{k+1,k}^{k}\right)&H_{k+1,k+1}^{k}\end{array}\right]}_{\coloneqq P}\underbrace{\left[\begin{array}[]{c}\sigma_{1}^{k}(t)I\\ \sigma_{2}^{k}(t)I\end{array}\right]}_{\coloneqq\Sigma}\succeq 0.$ (39)
The condition (39) can be achieved by satisfying $P\in\mathcal{C}$ where
$\mathcal{C}\coloneqq\left\{P\bigg{|}\forall\sigma_{1},\forall\sigma_{2}\in\mathbb{R}_{+},\Sigma=[\sigma_{1}I,\sigma_{2}I]^{\top},\Sigma^{\top}P\Sigma\succeq 0\right\}.$
It is worth noting that if $n_{H}=1$, the set $\mathcal{C}$ is equal to the set of $2\times 2$ copositive matrices [20]. With this observation, we are ready to derive sufficient LMI conditions ensuring $P\in\mathcal{C}$ by adapting techniques from the literature on matrix copositivity [12, 13]. It is worth noting that we formulate these conditions specifically for our problem so that we can obtain a finite set of LMIs that guarantee the invariance of the funnel.
Lemma 4.
Suppose that the following holds:
$H_{k,k}^{k}$ $\succeq 0,\quad H_{k+1,k+1}^{k}\succeq 0,$ (40a)
$H_{k,k+1}^{k}$ $+H_{k+1,k}^{k}\succeq 0.$ (40b)
Then, (39) holds for all $t$ in $(t_{k},t_{k+1})$.
Proof.
Each block of $P$ is PD because of the hypothesis in the Lemma 4. Since each $\sigma_{i}^{k}(t)$ for $i\in\{1,2\}$ is nonnegative for all $t$ in $(t_{k},t_{k+1})$, we have (29). ∎
The similar results were discussed in [21, 3]. Here, we can further derive a less conservative LMI condition implying $P\in\mathcal{C}$ compared to (40).
Lemma 5.
Suppose that there exist $H_{i,j}^{k}$ and $X_{i,j}^{k}$ for all $i,j\in\{k,k+1\}$ such that
$\left[\begin{array}[]{cc}H_{k,k}^{k}-X_{k,k}^{k}&\star\\ \frac{1}{2}\left(H_{k,k+1}^{k}+H_{k+1,k}^{k}\right)-X_{k+1,k}^{k}&H_{k+1,k+1}^{k}-X_{k+1,k+1}^{k}\end{array}\right]$ $\succeq 0,$ (41c)
$X_{i,j}^{k}=(X_{j,i}^{k})^{\top}\succeq 0,\quad\forall i,j\in\{k,k+1\}.$ (41d)
Then, (39) holds for all $t$ in $(t_{k},t_{k+1})$.
Proof.
Notice that the matrix $P$ defined in (39) can be written as
$P$ $=P_{1}+P_{2},$
where $P_{1}$ and $P_{2}$ are given by
$P_{1}$ $=\left[\begin{array}[]{cc}H_{k,k}^{k}-X_{k,k}^{k}&\star\\ \frac{1}{2}\left(H_{k,k+1}^{k}+H_{k+1,k}^{k}\right)-X_{k+1,k}^{k}&H_{k+1,k+1}^{k}-X_{k+1,k+1}^{k}\end{array}\right],$
$P_{2}$ $=\left[\begin{array}[]{cc}X_{k,k}^{k}&\star\\ X_{k+1,k}^{k}&X_{k+1,k+1}^{k}\end{array}\right].$
We have $\Sigma^{\top}P_{1}\Sigma\succeq 0$ because of the hypothesis in Lemma 5. Also, since every block $X_{i,j}^{k}$ for all $i,j\in\{k,k+1\}$ in $P_{2}$ is assumed to be PD, we have $\Sigma^{\top}P_{2}\Sigma\succeq 0$. Hence, we can conclude $\Sigma^{\top}P\Sigma\succeq 0$. ∎
By solving either (40) or (41) for all $k$ in $\mathbb{Z}_{[0,N-1]}$, we can guarantee the satisfaction of the DLMI (26) for all $t\in[t_{0},t_{f}]$ except at each temporal node points $t=t_{k}$ for all $k$ in $\mathbb{Z}_{[0,N]}$.
Remark 1.
The matrix-valued function $Q(t)$ with the FOH (8) is not differentiable at each temporal node points $t=t_{k}$ for all $k$ in $\mathbb{Z}_{[0,N]}$. This does not compromise our objective of continuous-time invariance since it suffices for the Lyapunov condition (20) to hold for almost all $t$ in $[t_{0},t_{f}]$ according to Lemma 1.
It is worth noting that (41) is less conservative than (40), as every solution of (40) is a special case of (41). This follows from the fact that (41) with the following constraints
$H_{i,i}^{k}$ $=X_{i,i}^{k},\quad\forall i\in\{k,k+1\},$
$\frac{1}{2}\left(H_{k,k+1}^{k}+H_{k+1,k}^{k}\right)$ $=X_{k,k+1}^{k},$
is equivalent to (40). However, solving (41) requires to introduce additional variables $X\in\mathbb{R}^{n_{H}\times n_{H}}$ that are larger in dimension than the variables $Q$ and $Y$, making the solution of (41) more computationally expensive than that of (40).
3.2 Constraints
The proposed funnel synthesis algorithm aims to satisfy not only the invariance of the funnel but also the state and input constraints. We consider linear state and input constraints written as
$\mathcal{P}_{x}$ =$=\{x\mid(a_{i}^{h})^{\top}x\leq b_{i}^{h},\quad i=1,\ldots,m_{x}\},$ (42a)
$\mathcal{P}_{u}$ =$=\{x\mid(a_{j}^{g})^{\top}u\leq b_{j}^{g},\quad j=1,\ldots,m_{u}\}.$ (42b)
We aim to make the state and input funnels at each node point centered around the nominal trajectory remain inside the feasible region. This can be stated using the set inclusions as $\{\bar{x}_{k}\}\oplus\mathcal{E}_{Q}(t_{k})\subseteq\mathcal{P}_{x}$ and $\{\bar{u}_{k}\}\oplus\mathcal{E}_{KQK^{\top}}(t_{k})\subseteq\mathcal{P}_{u}$ for all $k\in\mathbb{Z}_{[0,N]}$ where $\bar{x}_{k}=\bar{x}(t_{k})$ and $\bar{u}_{k}=\bar{u}(t_{k})$. These conditions can be equivalently written by the following LMIs [5]:
00 0 $\preceq\left[\begin{array}[]{cc}\left(b_{i}^{h}-(a_{i}^{h})^{\top}\bar{x}_{k}\right)^{2}&(a_{i}^{h})^{\top}Q_{k}\\ Q_{k}a_{i}^{h}&Q_{k}\end{array}\right],i=\mathbb{Z}_{[1,m_{x}]},$ (43c)
00 0 $\preceq\left[\begin{array}[]{cc}\left(b_{j}^{g}-(a_{j}^{g})^{\top}\bar{u}_{k}\right)^{2}&(a_{j}^{g})^{\top}Y_{k}\\ Y_{k}^{\top}a_{j}^{g}&Q_{k}\end{array}\right],j=\mathbb{Z}_{[1,m_{u}]}.$ (43f)
3.3 Problem formulation
The proposed funnel synthesis algorithm solves the following SDP problem:
[…]
(43)$,\forall k\in\mathbb{Z}_{[0,N]},$ (44c)
$0\prec Q_{k}\preceq Q_{max},\forall k\in\mathbb{Z}_{[0,N]}.$ (44d)
The cost function $J$ is assumed to be convex in $Q_{k}$. The matrix $Q_{max}\in\mathbb{S}_{++}^{n_{x}}$ is introduced to prohibit the funnel being arbitrarily large. By imposing $0\prec Q_{k}$ for each $k\in\mathbb{Z}_{[0,N]}$, the FOH interpolation (8) ensures $0\prec Q(t)$ for all $t\in[t_{0},t_{f}]$.
Theorem 1.
Let $Q(t)$ and $Y(t)$ be obtained from the solution of (44) with the FOH interpolation (8). Define the feedback gain $K(t)=Y(t)Q(t)^{-1}$. Then, the funnel $\mathcal{F}(t)$ defined in (19) is invariant for all $t\in[t_{0},t_{f}]$.
Proof.
By Lemma 4 and 5, satisfying (44b) implies that the DLMI (26) holds for all $t\in[t_{0},t_{f}]$ except at each temporal node point $t=t_{k}$ for all $k\in\mathbb{Z}_{[0,N]}$. It follows from Lemma 3 that the Lyapunov condition (20) holds almost everywhere. This implies the invariance of the state funnel $\mathcal{E}_{Q}$ followed from Lemma 1. Last, it follows from Lemma 2 that the funnel $\mathcal{F}(t)$ is invariant. ∎
Remark 2.
The feasibility condition in (43) enforces the constraints only at the temporal node points, which could result in violations between these node points. A possible remedy is to increase the number of node points. It is important to note that while the feasibility is only enforced at each node point, the invariance holds across the entire horizon $[t_{0},t_{f}]$ as established in Theorem 1.
3.4 Estimation of the constants via sampling
To solve the problem (44), it is necessary to obtain the local Lipschitz constant $(\gamma_{i})_{k}$ and the bound $\beta_{k}$ described in (10) for each $k\in\mathbb{Z}_{[0,N-1]}$. If the system is globally Lipschitz, one can directly use the global Lipschitz constant for all $k$ [22]. One effective way for estimating local $\gamma_{k}$ is to use sampling approaches [3], collecting samples of the triple $(\eta_{s},\xi_{s},w_{s})$ around the nominal trajectory, with $s$ denoting each sample. Specifically, this involves sampling the state deviation $\eta_{s}$ within the maximum state funnel $Q_{max}$, the input deviation $\xi_{s}$ within the constraint (42b), and the disturbance within its bound (2) for each subinterval $[t_{k},t_{k+1}]$. To obtain a less conservative constant, one might use the iterative procedure provided in [3] where the estimation of $\gamma_{k}$ and funnel computation are alternately repeated until the convergence. This paper does not consider this iterative process for simplicity as our proposed method can be incorporated with any way to estimate the Lipschitz constant. On the other hand, the value of $\Delta(t)$ defined in (6) is known and can thus be evaluated at each $t$. Hence, for each subinterval $[t_{k},t_{k+1}]$, we pick $N_{s}$ sample temporal points $t_{s}$ uniformly, and then $\beta_{k}$ can be determined through
$\beta_{k}=\max_{s=1,\ldots,N_{s}}\lVert\Delta(t_{s})\rVert_{2}.$
4 Numerical simulation
We consider the unicycle model given in (4) with $c_{1}=0.03$ and $c_{2}=0.05$. We divide a $15$ second time horizon into $N=20$ subintervals, starting at $t_{0}=0$ and ending at $t_{f}=15$. The cost function $J$ in (44) is set as $-\text{trace}(Q_{0})+\text{trace}(Q_{N})$ to maximize the funnel entry and minimize the funnel exit. The nominal trajectory illustrated in Fig. 2 starts at $(0,0)$, passes through $(4,8)$, and ends at $(8,0)$. A state constraint for avoiding an obstacle is taken into account, resulting in a nonconvex state constraint as shown in Fig. 2. This nonconvex constraint is linearized around the nominal trajectory to generate the polytopic constraints in the form of (42). The constraints for inputs are defined as follows: $0\leq u_{1}\leq 2$ (m/s) and $\lvert u_{2}\rvert\leq 1.5$ (rad/s). The constant $\lambda_{w}$ is chosen as $0.3$ and the parameter $Q_{max}$ for the maximum funnel size is set to $\text{diag}(2^{2},2^{2},(20\pi/180)^{2})$. To estimate $(\gamma_{i}){k}$ and $\beta_{k}$, we sample 100 triples of $(\eta_{s},\xi_{s},w_{s})$ and 20 samples of $\Delta(t_{s})$ for each subinterval $[t_{k},t_{k+1}]$. The Mosek solver is used to solve the SDP (44). The code for the simulation can be found in the following repository.
https://github.com/taewankim1/funnel-copositivity
Figure 2: (Top) The synthesized state funnel projected on $x$ ($x_{1}$) and $y$ ($x_{2}$) position coordinates. (Bottom) Time history of the state funnel projected on yaw angle ($x_{3}$) coordinate.
Figure 3: Time history of the synthesized input funnel projected on velocity command ($u_{1}$) and angular velocity command ($u_{2}$) coordinates, shown in the top and bottom figures, respectively.
Figure 4: (Top) The results of trajectories propagated from randomly selected samples within the funnel entry. (Bottom) Time history of Lyapunov function $V$, as defined in (15), for each trajectory sample.
The results of the state funnel and input funnel are provided in Fig. 2 and Fig. 3, respectively. It is clear that state and input constraints are satisfied at each node point. We conduct a comparison between two different conditions, (40) and (41), in terms of the resulting cost and computation time for synthesizing the funnel. The funnel computed using (41) has smaller cost, at -1.670, compared to -1.615 for the funnel using (40). The computational time for MOSEK to solve the SDP (44) with (40) is longer, at 8.34s, compared to 1.64s with (41), which is aligned with our expectation.
Finally, to test the invariant condition of the funnel, we take a total of 100 samples from the surface of the funnel entry $Q_{0}$. Then, these samples are propagated through the control law (21). During the propagation, we randomly choose the disturbance $w$ such that $\lVert w\rVert_{2}=1$. In the bottom figure of Fig. 4, the value of the Lyapunov function $V$ for each sample over the time horizon is illustrated. We can see that the values maintain below one, which means all trajectories remain in the funnel by the invariance property.
5 Conclusion
This paper presents a funnel synthesis algorithm for nonlinear systems under bounded disturbances. The proposed method can satisfy the invariance condition exactly over the finite time horizon. This is achieved by approximating the incremental system as an uncertain LPV system and deriving a finite number of LMIs that imply the DLMI resulting from the invariance condition. Simulation results show that the synthesized funnel satisfies both invariant and feasibility conditions.
Acknowledgement
This work was supported in part by the Air Force Office of Scientific Research under Grant FA9550-20-1-0053 and in part by the Office of Naval Research under Grant ONR N00014-20-1-2288.